Assume that an insulated box exits with two compartments A and B. Both compartments contain dry air and are separated by a wall that does not allow any heat through it but that is flexible enough to ensure equalization of pressure in both compartments. The initial conditions for both compartments are same T=273K, V=10^(-3)cubic meter, P=1013hpa. By means of electrical resistance, heat is supplied to the gas in A until its pressure becomes 10times initial pressure.Estimate
a) the final temperature in gas A
b) the work performed on the gas in B
c) the final temperature of the gas in A
d) the heat absorbed by the gas in A
1
Expert's answer
2019-09-16T04:05:26-0400

a) pV=nRT. n= pV/RT. T = pV/nR

n=1013 x 10/ 8.314 x 273 = 4.463 moles.

T= 1013 x 10 x 10 / 4.463 x 8.314 = 2730.06 K


b) if there are V= const, A= 0.


c) T= 1013 x 10 x 10 / 4.463 x 8.314 = 2730.06 K



d) ∆T = T1 - T2 = 2730.06 - 273 = 2,457.06217

Q = inR∆T/ 2= 3 x 4.463 x 8.314 x 2457.06217 / 2 = 136.755346 kJ.